Common stock valuelong dash—Variable growth 

Personal Finance Problem   Home Place​ Hotels, Inc., is entering into a​ 3-year remodeling and expansion project. The construction will have a limiting effect on earnings during that​ time, but when it is​ complete, it should allow the company to enjoy much improved growth in earnings and dividends. Last​ year, the company paid a dividend of ​$3.20 It expects zero growth in the next year. In years 2 and​ 3, 3%growth is​ expected, and in year​ 4, 15%growth. In year 5 and​ thereafter, growth should be a constant 8%per year. What is the maximum price per share that an investor who requires a return of 17​% should pay for Home Place Hotels common​ stock?
Answer #1

The maximum price is computed as shown below:

= Dividend in year 1 / (1 + required rate of return)1 + Dividend in year 2 / (1 + required rate of return)2 + Dividend in year 3 / (1 + required rate of return)3 + Dividend in year 4 / (1 + required rate of return)4 + 1 / (1 + required rate of return)4 [ ( Dividend in year 4 ( 1 + growth rate ) / ( required rate of return - growth rate ) ]

= $ 3.20 / 1.17 + ( $ 3.20 x 1.03 ) / 1.172 + ( $ 3.20 x 1.032 ) / 1.173 + ( $ 3.20 x 1.032 x 1.15 ) / 1.174 + 1 / 1.174 [ ( $ 3.20 x 1.032 x 1.15 x 1.08 ) / ( 0.17 - 0.08 ) ]

= $ 3.20 / 1.17 + $ 3.296 / 1.172 + $ 3.39488 / 1.173 + $ 3.904112 / 1.174 + $ 46.849344 / 1.174

= $ 34.35 Approximately

Feel free to ask in case of any query relating to this question
Answer #2

SOLUTION :


(Values in $) :


Last dividend, D0  = 3.20


No dividend growth next year.


Dividend next year, D1 = 3.20



Dividend growth in year 2 and 3 = 3.0% = 0.03


Dividend in year 2, D2 = 3.20*(1 + 0.03) = 3.296


Dividend in year 3, D3 = 3.296*(1 + 0.03) = 3.3949


Dividend growth in year 4 = 15% = 0.15


Dividend in year 4, D4 = 3.3949*(1 + 0.15) = 3.9041


Dividend growth in year 5 and in future = 8% = 0.08 (constant)


Dividend in year 5, D5 = 3.9041*(1 + 0.08) = 4.2164


So,


As per constant growth dividend model :


Required rate of return, r = 17% = 0.17

=> 1 + r = 1.17

Growth rate, g = 9% = 0.09 from year 5 onwards.


So, 


P4 (price at year 4 end) = D5/(r - g) = 4.2164/(0.17 - 0.08) = 46.85 


P0 (price today) 

= D1/(1+r) + D2/(1+r)^2 + D3/(1+r)^3 + (D4 + P4)/(1+r)^4

= 3.20/1.17+ 3.296/1.17^2 + 3.3949/1.17^3 + (4.2164+46.85)/1.17^4

= 34.51


Hence,


Maximum price that can be paid for 17% rate of return = $34.51 per share  (ANSWER).

answered by: Tulsiram Garg